OpenStudy (anonymous):

****- Find the rate of change between the two points: (6, 8) and (5, 6) where x is in days and y is in pounds.

OpenStudy (ipwnbunnies):

This is kinda like saying what is the slope of the line created by the two points. Recall slope formula: \[\frac{Y2-Y1}{X2-X1} = \frac{rise}{run}\]

OpenStudy (anonymous):

So 6^2-6^1 over 8^2-5^1 ?

OpenStudy (ipwnbunnies):

No, not 6 squared. I'm just distinguishing which points you should use. The difference between the 2 y-values divided by the difference between the 2 x-values.

OpenStudy (anonymous):

oh okay!

OpenStudy (ipwnbunnies):

What did you get? What math did you do?

OpenStudy (anonymous):

im in geometry

OpenStudy (ipwnbunnies):

You should know slope formula from Algebra 1 lol.

OpenStudy (ipwnbunnies):

\[\frac{Y2-Y1}{X2-X1} = \frac{8-6}{6-5} \frac{pounds}{day}\]